OpenStudy (anonymous):

I will fan and medal you. Solve each equation for the given variable. 5a + 14b = 23

OpenStudy (anonymous):

First of all, go steelers. Secondly you would make the a a zero so it would become 14b=23 and b becomes 23/14. then put in a zero for b and then it would become 5a=23, so a =23/5.

OpenStudy (anonymous):

does that help at all?

OpenStudy (anonymous):

a little i guess

OpenStudy (anonymous):

all you have to do is put in a zero for one of the variables then solve for the other. the way I solved for B was like this. \[5a+14b=23\] \[5(0)+14b=23\] \[14b=23\] \[\frac{ b=23 }{ 14 }\] \[b=\frac{ 23 }{ 14 }\]